A Bibliography of Supercomputing ’99

Nelson H. F. Beebe Center for Scientific Computing

University of Utah

Department of Mathematics, 322 INSCC

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu (Internet)
WWW URL: http://www.math.utah.edu/~beebe/

24 February 2000 Version 1.00

Abstract

This bibliography records articles presented at the Supercomputing ’99 conference.

Title word cross-reference

3 [AD99]. $7.3/Mflops [KFM99].

-D [AD99].

2 [AKL99]. 2000 [OSVW99].

5 [KFM99].

Accelerators [QRHD99]. Achieving [AGK^{+}99]. Adaptive [FSWB99, FL99, STY99]. ADR [KCS^{+}99].

Algorithms [BG99, CH99, GC99, TF99]. Alignments [CH99]. Analysis [VR99]. Application [AGK^{+}99, Joh99, OB99]. Applications [ABD^{+}99, CEV99, FSWB99, HKK^{+}99, KRC^{+}99, WSN99]. Approximate [YGS99]. Archipelago [MKY^{+}99]. Architectural [WMADC99]. Architecture [BLN^{+}99, CEV99, GC99, HKK^{+}99]. Architecture-cognizant [GC99]. area [GBST99]. Array [MMGL99]. ASCI [FJM^{+}99]. Astrophysical [KFM99]. Augmented [Fuc99]. Award [Nor99].

based [HKK^{+}99]. Beam [QRHD99]. Benchmarks [WMADC99]. benefit [VC99]. Biological [CH99]. BIP [GPT99]. BIP-SMP [GPT99]. Bit [ZZ99]. Bit-reversals [ZZ99]. Blue [FJM^{+}99].

Blue-Pacific [FJM^{+}99]. Body [KFM99]. Bottlenecks [SLT99]. Boulevard [ACM99]. Bounded [YGS99]. Bounded-error [YGS99].

Cache [SS99]. Cache-coherent [SS99]. Caches [RT99]. Calculation [HMM^{+}99]. Calculations [OSVW99]. Cameras [Fuc99]. Cell [QRHD99]. Center [ACM99]. CFD [AGK^{+}99]. Channel [PTC^{+}99]. Characteristics [WKKR99]. Chip [SHAS99]. Circulation [DH99, MKY^{+}99]. Climate [HJDW99, HHA99]. Cluster [GPT99, KEB^{+}99]. Clustering [MHHK99]. Clusters [GBST99, KRC^{+}99, SvL99]. Code [QRHD99]. cognizant [GC99]. coherent [SS99]. Collaborating [PRY99]. collaboration [Fuc99]. Collaboratory [PRY99]. Collective [MGF99]. Collectives [SvL99]. Combustion [PRY99]. Commodity [BLN^{+}99, GPT99, MDD^{+}99b]. Communication [BLN^{+}99, GGS99]. Community [HJDW99]. Comparisons [GGS99]. Compiler [ABD^{+}99]. Compiler-Supported [ABD^{+}99]. Compressible [MCC^{+}99]. Compression [YGS99]. Computation [HP99]. Computational [Fuc99, WBK^{+}99]. Computer [HMM^{+}99]. Conquer [GC99]. Constrained [KRC^{+}99, BG99]. Contrasts [GGS99]. Convention [ACM99]. Cost [VC99]. Cost-benefit [VC99].

D [AD99]. Data [SA99, DH99, FSWB99, FL99, GBST99, HKK^{+}99, KM99, MMGL99, YGS99]. Data-intensive [FSWB99, HKK^{+}99]. Datasets [KCS^{+}99]. DeepView [PTC^{+}99]. Design [MDD^{+}99a]. Diagnosis [KM99].

Diesel [PRY99]. dimensional [KCS^{+}99]. Direct [KEB^{+}99]. Distributed [FSWB99, PTC^{+}99]. DIVA [HKK^{+}99]. Divide [GC99]. DSM [SS99, SLT99]. Dynamic [KRC^{+}99, OB99, VR99]. Dynamics [QRHD99, WLMP99].

Efficient [BLN^{+}99, HP99, OB99, WSN99]. Electromagnetic [OSVW99]. Element [AD99, TF99]. Environment [MHHK99, SDN99]. Environments [LWF^{+}99]. error [YGS99]. Evaluating [ML99]. Evaluation [FJM^{+}99, HJDW99]. EveryWare [WBK^{+}99]. Execution [STY99]. Experiences [NvdP99].

Fact [KEB^{+}99].

Fast [STY99, WLMP99, ZZ99].

Fernbach [Nor99]. Fiction [KEB^{+}99].

Finite [AD99]. Flow [EDW99]. Fluid [EDW99]. Fortran90 [RY99]. Fujitsu [AKL99].

General [MKY^{+}99]. Generation [TFP99]. Generic [AKL99]. Geographic [MHHK99]. Global [MKY^{+}99]. Globus [MHHK99]. GRAPE [KFM99]. GRAPE-5 [KFM99]. Grid [LWF^{+}99, WBK^{+}99].

H [MRW99]. H-RMC [MRW99]. Hairpin [TFP99]. Hardware [SS99]. Heat [EDW99]. Hierarchical [SA99, YGS99]. High [AGK^{+}99, GPT99, HMM^{+}99, Joh99, MKY^{+}99, SS99, VC99, ZZ99, MCC^{+}99]. High-Performance [Joh99]. High-resolution [MKY^{+}99]. High-speed [HMM^{+}99]. Historical [KM99]. Hybrid [MRW99].

I/O [DH99, WSN99].

IBM [MCC^{+}99, SHAS99]. IBM-SP [MCC^{+}99]. ILU [HP99]. Imaging [MDD^{+}99b]. Immersive [TFP99]. Implementation [AKL99, EDW99]. Implementations [TF99, WLMP99]. Improved [CH99]. Improving [KM99, PH99]. Indonesian [MKY^{+}99]. Industrial [MDD^{+}99b]. Informatics [PTC^{+}99]. Informed [MGF99]. Input [MGF99]. Input/Output [MGF99]. Instruction [CEV99, MCFT99]. Instruction-level [MCFT99]. Integrated [MDD^{+}99a]. intensive [FSWB99, HKK^{+}99]. Interactive [RCLL99]. Internet [PRY99]. Irregular [HKK^{+}99]. Iterative [WKKR99].

Java [GGS99, MMGL99]. Java-MPI [GGS99]. Job [FJM^{+}99, LKK99]. Jr [ACM99].

Kernel [MRW99, AKL99]. King [ACM99].

Large [ABD^{+}99, EDW99, FL99, HMM^{+}99, KCS^{+}99, SvL99, WLMP99]. Large-scale [EDW99, HMM^{+}99, SvL99]. Level [RT99, GGS99, MCFT99, STY99]. Lighting [RCLL99]. Linear [QRHD99]. Linux [KEB^{+}99, MRW99]. Local [GBST99]. Locality [RT99]. Low [GGS99]. Low-level [GGS99]. Luther [ACM99].

Machines [STY99]. Management [STY99]. Managing [VR99]. Manufacturing [MDD^{+}99a]. Mapping [HKK^{+}99]. Martin [ACM99]. Matrix [CEV99, PH99]. Matrix-vector [PH99]. Memory [STY99, WKKR99]. Mesh [AGK^{+}99]. Message [AKL99, GPT99]. Metacomputing [MHHK99].

Methods [SA99, WKKR99, YGS99]. Microscopy [PTC^{+}99]. Mining [GBST99, MMGL99]. Model [DH99, HJDW99, MKY^{+}99]. MOE [HMM^{+}99]. Molecular [HMM^{+}99, WLMP99]. MOM [CEV99]. MPI [WSN99, AKL99, GGS99, STY99, SDN99, SvL99]. MPI-2 [AKL99]. MTA [ML99, OB99]. Multi [KCS^{+}99, MCFT99, RT99]. Multi-dimensional [KCS^{+}99]. Multi-Level [RT99]. Multi-threading [MCFT99]. Multicast [Joh99, MRW99]. Multigrid [AD99]. Multimedia [CEV99]. Multiphase [EDW99]. Multiple [LKK99]. Multiplication [PH99]. Multipole [WLMP99]. Multiprocessors [SS99, SLT99]. Multiresolution [FL99]. Multivariate [MHHK99].

N [KFM99]. N-Body [KFM99]. NAS [WMADC99]. Network [LWF^{+}99]. Nightmare [Fuc99]. Nirvana [Fuc99]. November [ACM99]. NREN [Joh99]. Numerical [KEB^{+}99, MKY^{+}99, TFP99].

O [DH99, WSN99]. Object [QRHD99]. Object-Oriented [QRHD99]. Ocean [DH99, MKY^{+}99]. Octrees [FL99]. Online [KM99]. Optimization [BG99, SvL99]. Optimizations [RT99]. Orbital [HMM^{+}99]. Oregon [ACM99]. Organization [DH99]. Oriented [QRHD99]. Origin [OSVW99]. Output [MGF99].

Pacific [FJM^{+}99]. Package [EDW99, MMGL99]. Papyrus [GBST99]. Parallel [AD99, ABD^{+}99, BG99, CH99, DH99, EDW99, FJM^{+}99, FL99, HJDW99, HMM^{+}99, HP99, MMGL99, QRHD99, WSN99, WMADC99].

Parallelism [MCFT99]. Parallelization [OB99, RCLL99]. Parallelizing [RY99]. Particle [QRHD99, YGS99]. Particle-in-Cell [QRHD99]. Passing [AKL99, GPT99]. PC [KEB^{+}99]. PC/Linux [KEB^{+}99]. PDE [BG99]. PDE-constrained [BG99]. Performance [AGK^{+}99, FSWB99, GPT99, GGS99, HJDW99, Joh99, KM99, LWF^{+}99, NvdP99, PH99, SS99, VC99, VR99, ZZ99]. Personal [HHA99]. PIM [HKK^{+}99]. PIM-based [HKK^{+}99]. Pinpointing [SLT99]. Portland [ACM99]. Preconditioners [HP99]. Prediction [FSWB99]. Predictive [VC99]. Prefetching [MGF99, VC99]. Preprocessor [RY99]. Presence [LKK99]. Prism [SDN99]. Problems [AD99]. Processors [MCFT99]. Programmable [RY99]. Programming [SDN99]. Programs [ML99]. Projection [VR99]. Projectors [Fuc99]. Protocol [MRW99]. Prototype [Joh99, NvdP99]. purpose [HMM^{+}99]. Pursuit [VR99].

Quantifying [SLT99]. Querying [KCS^{+}99].

Radiance [RCLL99]. Real [RCLL99]. Real-time [RCLL99]. Reality [Fuc99]. Recipient [Nor99]. Reliable [MRW99]. Requests [MGF99]. Requirements [LKK99, WMADC99]. Research [HHA99]. Researchers [PRY99]. Resolution [MCC^{+}99, MKY^{+}99]. Resource [LKK99]. reversals [ZZ99]. RMC [MRW99]. RS [SHAS99]. RS/6000 [SHAS99]. Running [WBK^{+}99].

SC’99 [ACM99]. Scal [SLT99]. Scal-Tool [SLT99]. Scalability [SLT99, WMADC99]. Scalable [OSVW99].

Scale [WLMP99, EDW99, HMM^{+}99, SvL99]. Scattering [OSVW99]. Scheduling [FJM^{+}99, KRC^{+}99, LKK99, SB99]. Scheme [VC99]. Seas [Fuc99]. Seismic [MDD^{+}99b]. Sequential [CH99]. Set [CEV99]. Sets [FL99]. SGI [OSVW99]. Shared [STY99]. Sidney [Nor99]. SIMD [CEV99]. Simulated [MKY^{+}99]. Simulation [ABD^{+}99, KEB^{+}99, KFM99, MCC^{+}99, QRHD99, TFP99]. Simulations [EDW99, WLMP99]. Simultaneous [MCFT99]. SMP [GPT99]. SMPs [GPT99, SvL99]. Software [EDW99]. Solver [AD99]. Sorting [SS99]. SP [MCC^{+}99, SHAS99]. Sparse [PH99]. Special [HMM^{+}99]. Special-purpose [HMM^{+}99]. Spectral [TF99]. speed [HMM^{+}99]. SPMD [ML99]. SQP [BG99]. Statistical [VR99]. Stochastic [SB99]. String [CH99]. Structure [SA99]. Sun [NvdP99, WSN99]. Super [GBST99]. Super-clusters [GBST99]. Supercomputer [HHA99]. Supercomputers [BLN^{+}99, MDD^{+}99b]. Support [SDN99]. Supported [ABD^{+}99]. Sustained [AGK^{+}99]. Switch [SHAS99]. System [GBST99, MCC^{+}99]. Systems [SHAS99, ZZ99].

Tele [Fuc99]. Tele-collaboration [Fuc99]. Tera [ML99, OB99]. Terascale [TF99]. Testbed [Joh99]. Thread [MCFT99, STY99]. Thread-level [MCFT99]. threading [MCFT99]. time [RCLL99]. Titanium [ML99]. Tool [LWF^{+}99, SLT99]. TOUGH2 [EDW99]. Tree [CH99]. Treecode [KFM99]. Tuning [HJDW99]. Tunnel [MKY^{+}99]. Turbulence [KEB^{+}99, MCC^{+}99]. Two [STY99]. Two-level [STY99].

Unifying [SA99].

Unstructured [AD99, AGK^{+}99, OB99]. Use [KM99]. Using [FL99, Joh99, MHHK99, MMGL99].

vector [PH99]. Very [ABD^{+}99, KCS^{+}99, MCC^{+}99]. Visualization [FL99, RCLL99, TFP99]. Vortex [TFP99]. vs [MCFT99].

Walking [CH99]. Walkthroughs [RCLL99]. Wide [GBST99]. Wide-area [GBST99]. WildFire [NvdP99]. Wind [MKY^{+}99].

References

Adve:1999:CSS

[ABD^{+}99] Vikram Adve, Rajive Bagrodia, Ewa Deelman, Thomas Phan, and Rizos Sakellariou. Compiler-supported simulation of very large parallel applications. In ACM [ACM99], page ??

ACM:1999:SPO

[ACM99] ACM, editor. SC’99: Oregon Convention Center 777 NE Martin Luther King Jr. Boulevard, Portland, Oregon, November 11–18, 1999. ACM Press and IEEE Computer Society Press, New York, NY 10036, USA and 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1999.

Adams:1999:PMS

[AD99] Mark Adams and Jim Demmel.

Parallel multigrid solver for 3-D unstructured finite element problems. In ACM [ACM99], page ??

Anderson:1999:AHS

[AGK^{+}99] Kyle Anderson, William Gropp, Dinesh Kaushik, David Keyes, and Barry Smith. Achieving high sustained performance in an unstructured mesh CFD application. In ACM [ACM99], page ??

Asai:1999:MIF

[AKL99] Noboru Asai, Thomas Kentemich, and Pierre Lagier. MPI-2 implementation on a Fujitsu Generic Message Passing Kernel. In ACM [ACM99], page ??

Biros:1999:PSA

[BG99] George Biros and Omar Ghattas. Parallel SQP algorithms for PDE-constrained optimization. In ACM [ACM99], page ??

Brauss:1999:ECA

[BLN^{+}99] Stephan Brauss, Martin Lienhard, Josef Nemecek, Anton Gunzinger, Martin Frey, Andreas Huber, Patrick Mueller, Martin Naef, and Roland Paul. An efficient communication architecture for commodity supercomputers. In ACM [ACM99], page ??

Corbal:1999:MMS

[CEV99] Jesus Corbal, Roger Espasa, and Mateo Valero. MOM: a matrix SIMD instruction set architecture for multimedia applications. In ACM [ACM99], page ??

Cull:1999:IPS

[CH99] Paul Cull and Tai-Ching Hsu.

Improved parallel and sequential walking tree algorithms for biological string alignments. In ACM [ACM99], page ??

Ding:1999:DOP

[DH99] Chris H. Q. Ding and Yun He. Data organization and I/O in a parallel ocean circulation model. In ACM [ACM99], page ??

Elmroth:1999:PIT

[EDW99] Erik Elmroth, Chris Ding, and Yu-Shu Wu. A parallel implementation of the TOUGH2 software package for large-scale multiphase fluid and heat flow simulations. In ACM [ACM99], page ??

Franke:1999:EPJ

[FJM^{+}99] Hubertus Franke, Joefon Jann, Jose Moreira, Pratap Pattnaik, and Morris Jette. An evaluation of parallel job scheduling for ASCI Blue-Pacific. In ACM [ACM99], page ??

Freitag:1999:AMV

[FL99] Lori A. Freitag and Raymond M. Loy. Adaptive, multiresolution visualization of large data sets using parallel octrees. In ACM [ACM99], page ??

Faerman:1999:APP

[FSWB99] Marcio Faerman, Alan Su, Richard Wolski, and Francine Berman. Adaptive performance prediction for distributed dataintensive applications. In ACM [ACM99], page ??

Fuchs:1999:ART

[Fuc99] Henry Fuchs. Augmented reality and tele-collaboration with seas of cameras and projectors: Computational nightmare or nirvana? In ACM [ACM99], page ?? Invited talk.

Grossman:1999:PSD

[GBST99] Robert L. Grossman, Stuart M. Bailey, Harinath Sivakumar, and Andrei L. Turinsky. Papyrus: A system for data mining over local and wide-area clusters and super-clusters. In ACM [ACM99], page ??

Gatlin:1999:ACD

[GC99] Kang Su Gatlin and Larry Carter. Architecture-cognizant divide and conquer algorithms. In ACM [ACM99], page ??

Getov:1999:MJM

[GGS99] Vladimir Getov, Paul Gray, and Vaidy Sunderam. MPI and Java-MPI: Contrasts and comparisons of low-level communication performance. In ACM [ACM99], page ??

Geoffray:1999:BSH

[GPT99] Patrick Geoffray, Loic Prylli, and Bernard Tourancheau. BIP-SMP: High performance message passing over a cluster of commodity SMPs. In ACM [ACM99], page ??

Hoe:1999:PSC

[HHA99] James C. Hoe, Chris Hill, and Alistar Adcroft. A personal supercomputer for climate research. In ACM [ACM99], page ??

Hammond:1999:PTE

[HJDW99] Steve Hammond, Rodney James, John B. Drake, and Patrick H. Worley. Performance tuning and evaluation of a parallel community climate model. In ACM [ACM99], page ??

Hall:1999:MIA

[HKK^{+}99] Mary Hall, Peter Kogge, Jeff Koller, Pedro Diniz, Jacqueline Chame, Jeff Draper, Jeff La-Coss, John Granacki, Apoorv Srivastava, William Athas, Jay Brockman, Vincent Freeh, Joonseok Park, and Jaewook Shin. Mapping irregular applications to DIVA, A PIM-based data-intensive architecture. In ACM [ACM99], page ??

Hashimoto:1999:MSP

[HMM^{+}99] Koji Hashimoto, Kazuaki Murakami, Nobuaki Miyakawa, Umpei Nagashima Hiroto Tomita, Koji Inoue, Katsuhiko Metsugi, Shinjiro Inabata, So Yamada, Hajime Takashima, Kunihiro Kitamura, Shigeru Obara, Takashi Amisaki, and Kazutoshi Tanabe. MOE: A specialpurpose parallel computer for high-speed, large-scale molecular orbital calculation. In ACM [ACM99], page ??

Hysom:1999:EPC

[HP99] David Hysom and Alex Pothen. Efficient parallel computation of ILU preconditioners. In ACM [ACM99], page ??

Johnson:1999:UNT

[Joh99] Marjory J. Johnson. Using the NREN Testbed to prototype a high-performance multicast application. In ACM [ACM99], page ??

Kurc:1999:QVL

[KCS^{+}99] Tahsin Kurc, Chialin Chang, Alan Sussman, Joel Saltz, and Renato Ferreira. Querying very large multi-dimensional datasets in ADR. In ACM [ACM99], page ??

Karamanos:1999:DNS

[KEB^{+}99] George-Sosei Karamanos, Constantinos Evangelinos, Richard C. Boes, Robert M. Kirby, and George E. Karniadakis. Direct numerical simulation of turbulence with a PC/Linux cluster: Fact or fiction? In ACM [ACM99], page ??

Kawai:1999:MAB

[KFM99] Atsushi Kawai, Toshiyuki Fukushige, and Junichiro Makino. $7.3/mflops astrophysical Nbody simulation with treecode on GRAPE-5. In ACM [ACM99], page ??

Karavanic:1999:IOP

[KM99] Karen Karavanic and Barton Miller. Improving online performance diagnosis by the use of historical performance data. In ACM [ACM99], page ??

Knobe:1999:SCD

[KRC^{+}99] Kathleen Knobe, James M.

Rehg, Arun Chauhan, Rishiyur S. Nikhil, and Umakishore Ramachandran. Scheduling constrained dynamic applications on clusters. In ACM [ACM99], page ??

Leinberger:1999:JSP

[LKK99] William Leinberger, George Karypis, and Vipin Kumar. Job scheduling in the presence of multiple resource requirements. In ACM [ACM99], page ??

Lee:1999:NPT

[LWF^{+}99] Craig A. Lee, Rich Wolski, Ian Foster, Carl Kesselman, and James Stepanek. A network performance tool for grid environments. In ACM [ACM99], page ??

Mirin:1999:VHR

[MCC^{+}99] A. A. Mirin, R. H. Cohen, B. C.

Curtis, W. P. Dannevik, A. M. Dimits, M. A. Duchaineau, D. E. Eliason, D. R. Schikore, S. E. Anderson, D. H. Porter, P. R. Woodward, L. J. Shieh, and S. W. White. Very high resolution simulation of compressible turbulence on the IBM-SP system. In ACM [ACM99], page ??

Mitchell:1999:ILP

[MCFT99] Nicholas Mitchell, Larry Carter, Jeanne Ferrante, and Dean Tullsen. Instruction-level parallelism vs. thread-level parallelism on simultaneous multithreading processors. In ACM [ACM99], page ??

Moran:1999:IMD

[MDD^{+}99a] David Moran, Gary Ditlow, Daria Dooling, Ralph Williams, Paul Smith, Tom Wilkins, Richard Moore, and Anshul Gupta. Integrated manufacturing and design. In ACM [ACM99], page ??

Morton:1999:ISI

[MDD^{+}99b] Scott A. Morton, Jeffrey R. Davis, Harry L. Duffey, Gary L. Donathan, Vic Forsyth, and Steven N. Checkles. Industrial seismic imaging on commodity supercomputers. In ACM [ACM99], page ??

Madhyastha:1999:IPC

[MGF99] Tara M. Madhyastha, Garth A. Gibson, and Christos Faloutsos. Informed prefetching of collective input/output requests. In ACM [ACM99], page ??

Mahinthakumar:1999:MGC

[MHHK99] G. (Kumar) Mahinthakumar, Forrest M. Hoffman, William W. Hargrove, and Nicholas T. Karonis. Multivariate geographic clustering in a metacomputing environment using Globus. In ACM [ACM99], page ??

Masumoto:1999:SCI

[MKY^{+}99] Masumoto, Takashi Kagimoto, Masahiro Yoshida, Masahiro Fukuda, Naoki Hirose, and Toshio Yamagata. Simulated circulation in the Indonesian Archipelago from a highresolution global ocean general circulation model on the numerical wind tunnel. In ACM [ACM99], page ??

Miyamoto:1999:ETS

[ML99] Carleton Miyamoto and Chang Lin. Evaluating Titanium SPMD programs on the Tera MTA. In ACM [ACM99], page ??

Moreira:1999:PDM

[MMGL99] Jose Moreira, Sam Midkiff, Manish Gupta, and Rick Lawrence. Parallel data mining using the array package for Java. In ACM [ACM99], page ??

McKinley:1999:HRH

[MRW99] Philip K. McKinley, Ravi T.

Rao, and Robin F. Wright. H-RMC: A hybrid reliable multicast protocol for the Linux kernel. In ACM [ACM99], page ??

Norman:1999:SFA

[Nor99] Michael Norman. Sidney Fernbach award recipient. In ACM [ACM99], page ??

Noordergraaf:1999:PES

[NvdP99] Lisa Noordergraaf and Ruud van der Pas. Performance experiences on Sun’s WildFire prototype. In ACM [ACM99], page ??

Oliker:1999:EPD

[OB99] Leonid Oliker and Rupak Biswas. Efficient parallelization of a dynamic unstructured application on the Tera MTA. In ACM [ACM99], page ??

Ottusch:1999:SES

[OSVW99] John J. Ottusch, Mark A. Stalzer, John L. Visher, and Stephen M. Wandzura. Scalable electromagnetic scattering calculations on the SGI Origin 2000. In ACM [ACM99], page ??

Pinar:1999:IPS

[PH99] Ali Pinar and Michael T. Heath. Improving performance of sparse matrix-vector multiplication. In ACM [ACM99], page ??

Pancerella:1999:DCC

[PRY99] Carmen M. Pancerella, Larry Rahn, and Christine Yang. The Diesel Combustion Collaboratory: Combustion researchers collaborating over the Internet. In ACM [ACM99], page ??

Parvin:1999:DCD

[PTC^{+}99] Bahram Parvin, John Taylor, Ge Cong, Michael O’Keefe, and Mary-Helen Barcellos-Hoff. DeepView: A channel for distributed microscopy and informatics. In ACM [ACM99], page ??

Qiang:1999:OOP

[QRHD99] Ji Qiang, Robert Ryne, Salman Habib, and Viktor Decyk. An object-oriented parallel particle-in-cell code for beam dynamics simulation in linear accelerators. In ACM [ACM99], page ??

Robertson:1999:PRR

[RCLL99] David Robertson, Kevin Campbell, Stephen Lau, and Terry Ligocki. Parallelization of radiance for real-time interactive lighting visualization walkthroughs. In ACM [ACM99], page ??

Rivera:1999:LOM

[RT99] Gabriel Rivera and Chau-Wen Tseng. Locality optimizations for multi-level caches. In ACM [ACM99], page ??

Rosing:1999:PPP

[RY99] Matthew Rosing and Steve B.

Yabusaki. A programmable preprocessor for parallelizing Fortran90. In ACM [ACM99], page ??

Aluru:1999:UDS

[SA99] Fatih Sevilgen, Sr.inivas Aluru.

A unifying data structure for hierarchical methods. In ACM [ACM99], page ??

Schopf:1999:SS

[SB99] Jennifer M. Schopf and Francine Berman. Stochastic scheduling. In ACM [ACM99], page ??

Sistare:1999:MSP

[SDN99] Steve Sistare, Erica Dorenkamp, and Nick Nevin. MPI support in the Prism programming environment. In ACM [ACM99], page ??

Stunkel:1999:NSC

[SHAS99] Craig B. Stunkel, Jay Herring, Bulent Abali, and Rajeev Sivaram. A new switch chip for IBM RS/6000 SP systems. In ACM [ACM99], page ??

Solihin:1999:STP

[SLT99] Yan Solihin, Vinh Lam, and Josep Torrellas. Scal-Tool: Pinpointing and quantifying scalability bottlenecks in DSM multiprocessors. In ACM [ACM99], page ??

Shan:1999:HPS

[SS99] Hongzhang Shan and Jaswinder P. Singh. High performance sorting on hardware cache-coherent DSM multiprocessors. In ACM [ACM99], page ??

Shen:1999:ATL

[STY99] Kai Shen, Hong Tang, and Tao Yang. Adaptive twolevel thread management for fast MPI execution on shared memory machines. In ACM [ACM99], page ??

Sistare:1999:OMC

[SvL99] Steve Sistare, Rolf vandeVaart, and Eugene Loh. Optimization of MPI collectives on clusters of large-scale SMPs. In ACM [ACM99], page ??

Tufo:1999:TSE

[TF99] H. M. Tufo and P. F. Fischer.

Terascale spectral element algorithms and implementations. In ACM [ACM99], page ??

Tufo:1999:NSI

[TFP99] Henry Tufo, Paul Fischer, and Mike Papka. Numerical simulation and immersive visualization of hairpin vortex generation. In ACM [ACM99], page ??

Vellanki:1999:CBS

[VC99] Vivekanand Vellanki and Ann Chervenak. A cost-benefit scheme for high performance predictive prefetching. In ACM [ACM99], page ??

Vetter:1999:MPA

[VR99] Jeffrey Vetter and Daniel Reed.

Managing performance analysis with dynamic statistical projection pursuit. In ACM [ACM99], page ??

Wolski:1999:REC

[WBK^{+}99] Rich Wolski, John Brevik, Chandra Krintz, Graziano Obertelli, Neil Spring, and Alan Su. Running EveryWare on the computational grid. In ACM [ACM99], page ??

Weiss:1999:MCI

[WKKR99] Christian Weiss, Wolfgang Karl, Markus Kowarschik, and Ulrich Rüde. Memory characteristics of iterative methods. In ACM [ACM99], page ??

Wang:1999:LSM

[WLMP99] Zhiqiang Wang, James Lupo, Alan McKenney, and Ruth Pachter. Large scale molecular dynamics simulations with fast multipole implementations. In ACM [ACM99], page ??

Wong:1999:ARS

[WMADC99] Frederick C. Wong, Richard P. Martin, Remzi H. Arpaci-Dusseau, and David E. Culler. Architectural requirements and scalability of the NAS Parallel Benchmarks. In ACM [ACM99], page ??

Wisniewski:1999:SME

[WSN99] Len Wisniewski, Brad Smisloff, and Nils Nieuwejaar. Sun MPI I/O: Efficient I/O for parallel applications. In ACM [ACM99], page ??

Yang:1999:BEC

[YGS99] Dow-Yung Yang, Ananth Y. Grama, and Vivek Sarin. Bounded-error compression of particle data from hierarchical approximate methods. In ACM [ACM99], page ??

Zhang:1999:FBR

[ZZ99] Zhao Zhang and Xiaodong Zhang. Fast bit-reversals on high performance systems. In ACM [ACM99], page ??